OpenStudy (anonymous):

Find the equation of the line perpendicular to the line 3x=y-5 and passing thruogh the point (7, 21)

OpenStudy (amistre64):

perp lines swap the coeffs of x and y then changes the sign of one of them

OpenStudy (amistre64):

3x = y - 5 ; perps to -x = 3y - 5 ; but the 5 is useless, AND this is not in any particular line format :)

OpenStudy (amistre64):

drop the f and calibrate with the given point

OpenStudy (amistre64):

f means 5 for some reason :)

OpenStudy (amistre64):

-x = 3y - 5 -x = 3y - n -(7) = 3(21) - n n = 63 + 7 n = 70 -x = 3y + 70 ...

OpenStudy (amistre64):

format it to whatever line format they will accept

OpenStudy (anonymous):

y=-1/3(x)+70 use thi method y=mx+c sub:x and y 21=-1/3(7)+c c=70 since the m of the original equation is 3 this means for perpendicular line will be -1/3